OpenStudy (anonymous):

Darla is building a new desk. To make sure she had made a square corner, she measures 5ft from the corner along one edge and 3ft from the corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle? Diagonal ? ft

OpenStudy (amistre64):

darla needs to learn carpentry if shes gonna try to make a desk

OpenStudy (saifoo.khan):

5^2+8^2=D^2 25+64=d^2 89=d^2 Sqrt(89)=Sqrt(d^2)=9.43ft

OpenStudy (amistre64):

tell her to measure 4 ft and 3ft and if the diag is 5ft its square enough

OpenStudy (saifoo.khan):

lol

OpenStudy (amistre64):

then, if she wants she can do these measurements and find out lol

OpenStudy (saifoo.khan):

haha!